Starting at t = 0 s , a horizontal net force F⃗ =( 0.290 N/s )ti^+(-0.440 N/s2 )t2j^ is applied to a box that has an initial momentum p⃗ = ( -2.85 kg⋅m/s )i^+( 4.00 kg⋅m/s )j^ . Part A What is the momentum of the box at t = 1.90 s ? Enter the x and y components of the momentum separated by a comma.
Answer #1

F= dp/dt

p = \intFdt

=> pf - pi = \int[( 0.290 N/s )t i+(-0.440 N/s2 )t2 j] dt

=> pf - (( -2.85 kg⋅m/s ) i+( 4.00 kg⋅m/s ) j) = [( 0.290 N/s )t2 /2 i+(-0.440 N/s2 )t3 /3 j]

For t= 1.90 s

=> pf - (( -2.85) i+( 4.00) j) = [( 0.290* (1.90)2 /2 i+(-0.440)*(1.90)3 /3 j]

=> pf = [( 0.290* (1.90)2 /2 i+(-0.440)*(1.90)3 /3 j] + (( -2.85) i+( 4.00) j

=> pf = -2.33 kg⋅m/s i + 2.99 kg⋅m/s j

s0, pf = -2.33 , 2.99